Answer is 1,000rpm
Problem 2 A fan with rotation speed of 950 rpm is able to manage a static pressure of 5 w.c. at sea level. If the same static pressure needs to be managed by the fan at an altitude of 2811 ft, what must be the new rotational speed?
Answer #1

GIVEN :

N1 = 950 rpm, P1 = P2 = 5" w.c, D1 = 1312 ft, D2 = 2811 ft

To find : N2

SOLN : The second fan law relates the fan total pressure or fan static pressure to the fan rotational speed: Total or static pressure, P, is proportional to the square of the fan rotational speed, N

(P1/P2) = (D1/D2)2 (N1/N2)2 (?1/?2)

(5/5) = (1312/2811)2 x (950/N2)2 x 5.08 (density ratio = 5.08 for the given conditions)

On calculating we get, N2 = 1000 rpm